Give first the step you will use to separate the variable and then solve the equation:
8y = 36

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Now we have to divide both sides of the equation by 8,
Then we get,
= 8y/8 = 36/8
= x = 9/4

Thanks